# Analysis of Clustering and Degree Index in Random Graphs and Complex Networks

### v1

@@ Abstract
The purpose of this paper is to analyze the degree index and clustering index in random graphs. The degree index in our setup is a certain measure of degree irregularity whose basic properties are well studied in the literature, and the corresponding theoretical analysis in a random graph setup turns out to be tractable. On the other hand, the clustering index, based on a similar reasoning, is first introduced in this manuscript. Computing exact expressions for the expected clustering index turns out to be more challenging even in the case of Erdős-Rényi graphs, and our results are on obtaining relevant upper bounds. These are also complemented with observations based on Monte Carlo simulations. Besides the Erdős-Rényi case, we also do simulation-based analysis for random regular graphs, the Barabási-Albert model and the Watts-Strogatz model.
@@ 1 Introduction
Since the introduction of Erdős-Rényi graphs around 1960 ([7], [8]), random graphs gained substantial importance in both applied and theoretical fields. Meanwhile, various statistics of interest for the underlying networks (or, graphs) have been studied, which include the average degree, clustering coefficient, efficiency, and modularity, among several others. See, for example, [6], [12], [14] and [20] for relevant definitions and certain applications. Due to the development of graph-based machine learning algorithms in recent years, these statistics have been even more popular, especially in classification problems [21]. The purpose of this paper is to study two other types of graph characteristics in random graphs.
In order to describe the two characteristics we study, let $\mathcal{G}=(V,E)$ be a graph where $V=\{1,2,\ldots,n\}$ for some $n\in\mathbb{N}$. Let $N(k)$ be the set of neighbors of $k\in V$ in $\mathcal{G}$. We denote the degree of $k$ by $d_{k}$. With this notation, we define the $\alpha$-level degree index of $\mathcal{G}$ to be
@@
Here, $\alpha>0$ is a real number which is chosen to be in $\{1,2\}$ in this manuscript. Also let us note that $\sum_{i<j}$ will mean $\sum_{1\leq i<j\leq n}$ below.
The second graph characteristic of interest for us will be the clustering index, which, as far as we know, is not studied previously in the literature. First, recall that for given $i\in V$, when $d_{i}\geq 2$, the local clustering coefficient of $i$ is defined to be
@@
In case where $d_{i}\in\{0,1\}$, we set $C(i)=0$. Now, similar to the degree index, for $\alpha>0$, we define the $\alpha$-level clustering index of $\mathcal{G}$ to be
@@
Here again $\alpha$ in our case is either 1 or 2. In the upcoming sections, we will analyze both $\mathrm{DI}_{\alpha}(G)$ and $\mathrm{CI}_{\alpha}(G)$ when $G$ is an Erdős-Rényi graph. We will complement these by doing observations on other random graph models.
The clustering index is a measure that provides certain information about the uniformity among the local neighborhood structure of $\mathcal{G}$ by focusing on the clustering. As already noted, although such a coefficient is quite natural to define, it is first introduced in this manuscript to the best of our knowledge. That said, the degree index defined above is previously studied in the literature in distinct contexts. Our primary goal in this paper is to analyze these two indices for the case of Erdős-Rényi graphs. For such random graphs, the degree index case turns out to be relatively straightforward, and precise formulas for the expected degree index are obtainable. On the other hand, the case of the clustering index is more involved, and we are not able to obtain exact formulas. However, we show that when $\mathcal{G}$ is an Erdős-Rényi graph with $n$ nodes and a fixed attachment probability $p$, the bounds
$\mathbb{E}[\mathrm{CI}_{1}(\mathcal{G})]\leq K_{1}n$
and
$\mathbb{E}[\mathrm{CI}_{2}(\mathcal{G})]\leq K_{2},$
hold for every $n$, where $K_{1},K_{2}$ are constants depending on $p$, but not on $n$. Our heuristic arguments yield that matching lower bounds are also true for both cases, but we have not verified these rigorously yet.
Let us now give some pointers for previous work on the degree index, and then include some discussions and references on the clustering coefficient. Noting that $\mathrm{DI}_{\alpha}(\mathcal{G})=0$ means that the graph under study is regular, this quantity is considered as a measure of irregularity of the graph [22]. An irregularity definition that has found significant use in the literature is the Albertson index introduced in [5], in which the corresponding definition is similar to ours, but the sum under consideration is over the edge set instead of the vertices. Although the irregularity used in the present manuscript is quite natural to consider, the special case $\mathrm{DI}_{1}(\mathcal{G})$ was relatively recently introduced in [1], where the authors call it the total irregularity of the graph. Also, a form of the special case of $\mathrm{DI}_{2}(\mathcal{G})$ in which the corresponding sum is over the edge set instead of the node was proposed in [2]. Aside from these, there are other notions of graph degree irregularity, see [3] for a partial list.
The clustering coefficient of a node $i\in V(\mathcal{G})$ is a measure of the likeliness of neighbors of $i$ to be neighbors among themselves [23]. The clustering coefficient is extensively studied in the literature, and several variations and generalizations are proposed; see [15], [16] and [17] for some exemplary work. The clustering coefficient is also studied in a random graph setup. For instance, [24] analyzes the clustering coefficient in Erdős-Rényi graphs and certain random regular graphs. [9] does an investigation in the case of a generalized small world model. On the other hand, [17] discusses a natural generalization of the standard clustering coefficient, and analyzes in both Erdős-Rényi and small world settings.
Here is a summary of the contributions of the current manuscript. The degree index we study is well known in the literature, but it is not analyzed for random graphs to the best of our knowledge. Here we are able to fully understand the expected degree index for $\alpha\in\{1,2\}$. We are not aware of any work on clustering index in the literature. After introducing this concept, we prove related results on the expected clustering index for Erdős-Rényi graphs, and in particular we obtain upper bounds for this quantity for $\alpha\in\{1,2\}$. Although the theoretical work is done for the Erdős-Rényi due to its simplicity, we complement this with a simulation study on other random graph models.
@@
The rest of the paper is organized as follows. Next section contains some basic observations on the clustering index, which can be easily adapted to the degree index. This section also includes a brief discussion on the comparison of the extremal cases of the degree index and the clustering index. Afterwards, in Section 3 we analyze the clustering index in Erdős-Rényi graphs, and provide upper bounds for the expected clustering index for both $\alpha=1$ and $2$. Later, in Section 4, a similar study is done for the degree index. Section 5 is then devoted to a simulation based study for the indices under study for random regular graphs, Barabási-Albert model and Watts-Strogatz model. Lastly, we conclude the paper in Section 6 with some discussions on possible future work.
@@ 2 Some basic observations on the clustering index
In this section we do some elementary observations on the clustering index. Towards this, we first analyze the extremal values of this index. Below, whenever it is clear from the context, we assume that the node set of the underlying graph is $V=\{1,2,\ldots,n\}$ for some $n\in\mathbb{N}$.
Proposition 2.1
For any graph $\mathcal{G}=(V,E)$ and for any $\alpha\geq 1$, we have
$0\leq\mathrm{CI}_{\alpha}(\mathcal{G})\leq\frac{n^{2}}{4}.$
Lower bound of the proposition is clear. The upper bound follows from the following elementary observation.
Proposition 2.2
Let $a_{1},a_{2},\ldots,a_{n}$ be real numbers over $[0,1]$. Then for any $\alpha\geq 1$, we have,
@@
The reason for the last equality is that for the terms containing $b_{k}$, it must be that $i\leq k-1$ and $j\geq k$. The number of such pairs is $(k-1)(n+1-k)$. The maximum value $(k-1)(n+1-k)$ can take is $\frac{n^{2}}{4}$. Therefore, $S\leq\frac{n^{2}}{4}\sum_{k=2}^{n}b_{k}=$ $\frac{n^{2}}{4}\left(a_{n}-a_{1}\right)\leq\frac{n^{2}}{4}$. $\square$
The following example discusses some elementary graphs for which the bounds stated in Proposition 2.1 are attained.
@@
Let $\alpha=1$. The lower bound in Proposition 2.1 is attained, for example, for the following graphs:
@@
As a simple example where the clustering index is maximized, let $m\geq 2$, $n=2m$, and consider the graph that is the graph union of an $m$-null set, and an $m$-clique. Here the only contribution to the sum defining $\mathrm{CI}$ is from the cases where a node from the $m$-clique, and another node from the $m$-null set are considered. Noting that the clustering coefficient of the node from the $m$-null set is zero by definition, the observation that $\mathrm{CI}$ takes the value $n^{2}/4$ follows immediately.
Another related example will be given Example 2.3 where extremal situations in clustering and degree indices are compared and discussed.
Noting that the observations in Proposition 2.1 can be easily adapted to the degree index, we continue with the following result which provides a way to express the clustering index in terms of the order statistics of the corresponding local clustering coefficients.
Proposition 2.3
@@
Proof. Following the proof of Proposition 2.2 with $a_{i}=C_{(i)}$ gives
@@
Let us also examine extremal cases in random graphs. Clearly, any random tree model provides a random graph model where the lower bound in Proposition 2.1 is achieved. The following provides a simple random model where the upper bound in Proposition 2.1 is achieved asymptotically.
@@
Begin with two empty sets of nodes $V_{1}$ and $V_{2}$ at time 0. At each time $t\in\mathbb{N}$, a newly generated node $v_{t}$ is inserted in $V_{1}$ or $V_{2}$ with probabilities $p$ and $1-p$, respectively. If $v_{t}$ is inserted in $V_{1}$, we do not attach it to any present nodes. If it is inserted in $V_{2}$, then we attach it to all vertices present in $V_{2}$. Then at time $n\in\mathbb{N}$, the vertices in $V_{1}$ form a null-set (or, an independent set), and the vertices in $V_{2}$ form a clique.
Letting $N$ be the number of vertices in $V_{1}$ at time $n$, $N$ is binomially distributed with parameters $n$ and $p$, and we have,
@@
Next we discuss examples towards a comparison of the clustering index and the degree index. In particular, we are interested in cases where the degree index is high and clustering coefficient is low, and vice versa.
@@
Consider a collection of disjoint polygons, i.e. a graph where each node is contained in a unique polygon. $C(i)=1$ if $i$ is contained in a triangle, and $C(i)=0$ if $i$ is contained in a polygon other than a triangle. The degree of any node in this graph is 2. Assuming we have $k$ vertices contained in some triangle and $n-k$ vertices contained in a polygon which is not a triangle, we have $\mathrm{CI}_{\alpha}=\sum\limits_{1\leq i<j\leq n}|C(i)-C(j)|^{\alpha}=nk-k^{2}$. If $n$ is even, then the maximum is attained when $k=n/2$, which yields $n^{2}/4$. Thus we have found a graph for which the degree index is zero, but the clustering index is $n^{2}/4$, the maximal possible value.
@@
Next, we also discuss an example where the clustering index is zero while the degree index $\sum\limits_{1\leq i<j\leq n}|d_{i}-d_{j}|$ is $\frac{n^{2}(n/2-3)}{4}$ (Proposition 2.2 can be used to show that the degree index can be at most $n^{3}/4$. So our example here is off by a factor of $1/2$, but still demonstrates that the degree index can be very large while the clustering index is zero.).
@@ 3 Clustering index of Erdős-Rényi graphs
In this subsection we will show that $\mathbb{E}[\mathrm{CI}_{1}(\mathcal{G})]$ and $\mathbb{E}[\mathrm{CI}_{2}(\mathcal{G})]$ are both sublinear. The result for the latter will later be improved, and in particular $\mathbb{E}[\mathrm{CI}_{2}(\mathcal{G})]$ will be shown to be bounded by a constant independent of $n$. We begin with the following proposition concerning the first two moments of the local clustering coefficient.
@@
$\mathbb{E}[C(i)]=\phi\mathbb{E}[C(i)\mid d_{i}\leq 1]+(1-\phi)\mathbb{E}[C(i)\mid d_{i}\geq 2].$
Now, by definition of $C(i)$, $\mathbb{E}[C(i)\mid d_{i}\leq 1]=0$. Also, $\mathbb{E}[C(i)\mid d_{i}\geq 2]=p$. Hence, we obtain
$\mathbb{E}[C(i)]=p(1-(1-p)^{n-1}-(n-1)p(1-p)^{n-2}).$
The assertion that $\mathbb{E}[C(i)]\geq p-\frac{D}{n^{2}}$ is now clear.
(ii) Note that, below $D$ denotes a constant independent of $n$ that does not necessarily have the same value in two appearances.
Let $A_{i}$ be the event that $|d_{i}-(n-1)p|<n^{2/3}$. Then by McDiarmid (or, Azuma-Hoeffding) inequality [18], $\mathbb{P}(\overline{A_{i}})\leq M_{1}e^{-M_{2}n^{1/3}}$ for some positive constants $M_{1},M_{2}$ independent of $n$. Using $A_{i}$, we now write,
@@
Now, when $A_{i}$ is true $(n-1)p-n^{2/3}\leq d_{i}\leq(n-1)p+n^{2/3}$. Keeping this in mind, and using the last relation we obtained, we get,
@@
Proof. Using Cauchy-Schwarz inequality, and the previous proposition,
@@
In the next subsection we will show that in case of Erdős-Rényi graphs for fixed $p$, $\mathbb{E}[\mathrm{CI}_{2}(\mathcal{G})]$ is indeed bounded by a constant that does not depend on $n$. That said, let us note that our simulation results further suggest that for Erdős-Rényi graphs, the growth of $\mathrm{CI}_{1}(\mathcal{G})$ behaves linearly in the number of nodes, and $\mathrm{CI}_{2}(\mathcal{G})$ converges to a positive constant as the number of nodes increases.

### v2

@@ Abstract
The purpose of this paper is to analyze the degree index and the clustering index in random graphs. The degree index in our setup is a certain measure of degree irregularity whose basic properties are well studied in the literature, and the corresponding theoretical analysis in a random graph setup turns out to be tractable. On the other hand, the clustering index, based on a similar reasoning, is first introduced in this manuscript. Computing exact expressions for the expected clustering index turns out to be more challenging even in the case of Erdős-Rényi graphs, and our results are on obtaining relevant upper bounds. These are also complemented with observations based on Monte Carlo simulations. Besides the Erdős-Rényi case, we also do simulation-based analysis for random regular graphs, the Barabási-Albert model and the Watts-Strogatz model.
Keywords: Random graphs, Erdős-Rényi graphs, clustering coefficient, clustering index, degree irregularity.
MSC2020 subject classifications: 05C80, 05C82, 05C07, 60C05.
@@ 1 Introduction
Since the introduction of Erdős-Rényi graphs around 1960 ([10], [11]), random graphs have gained substantial importance in both applied and theoretical fields. Meanwhile, various statistics of interest for the underlying networks (or graphs) have been studied, including the average degree, clustering coefficient, efficiency, and modularity, among several others. See, for example, [8], [13], [15] and [22] for relevant definitions and certain applications. Due to the development of graph-based machine learning algorithms in recent years, these statistics have become even more popular, especially in classification problems [23]. The purpose of this paper is to study two other types of graph characteristics in random graphs.
In order to describe the two characteristics that we study, let $\mathcal{G}=(V,E)$ be a graph where $V=\{1,2,\ldots,n\}$ for some $n\in\mathbb{N}$. Let $N(k)$ be the set of neighbors of $k\in V$ in $\mathcal{G}$. We denote the degree of $k$ by $d_{k}$. With this notation, we define the $\alpha$-level degree index of $\mathcal{G}$ to be
@@
Here, $\alpha>0$ is a real number, which is chosen to be in $\{1,2\}$ in this manuscript. Also, let us note that $\sum_{i<j}$ will mean $\sum_{1\leq i<j\leq n}$ below. Noting that $\mathrm{DI}_{\alpha}(\mathcal{G})=0$ means that the graph under study is regular, this quantity is considered as a measure of irregularity of the graph [24]. On the other hand, in networks with high disparity in node connectivity, the index $\mathrm{DI}_{\alpha}(\mathcal{G})$ would yield large values. The index overall provides a global measure of degree variability in the network and may, for example, help differentiate between an Erdős-Rényi graph and a Barabási-Albert network even if they have the same number of nodes and edges.
A definition of degree irregularity that has found significant use in the literature is the Albertson index introduced in [5], in which the corresponding definition is similar to ours, but the sum under consideration is over the edge set instead of the vertices. Although the irregularity used in the present manuscript is quite natural to consider, the special case $\mathrm{DI}_{1}(\mathcal{G})$ was relatively recently introduced in [1], where the authors call it the total irregularity of the graph. In addition, a form of the special case of $\mathrm{DI}_{2}(\mathcal{G})$, in which the corresponding sum is over the edge set instead of the vertex set, was proposed in [2]. Aside from these, there are other notions of graph degree irregularity; see [3] for a partial list and [6] for a broad look at the topic.
The second graph characteristic of interest for us will be the clustering index, which, as far as we know, is not studied previously in the literature. Before introducing this index, recall that for given $i\in V$, when $d_{i}\geq 2$, the local clustering coefficient of $i$ is defined to be
@@
In the case where $d_{i}\in\{0,1\}$, we set $C(i)=0$. In general, the clustering coefficient of a node $i\in V(\mathcal{G})$ is a measure of the likelihood of the neighbors of $i$ to be neighbors among themselves [25]. The clustering coefficient is extensively studied in the literature, and several variations and generalizations are proposed; see [16], [18] and [26] for some exemplary work. The clustering coefficient is also studied in a random graph setup. For instance, [17] analyzes the clustering coefficient in Erdős-Rényi graphs and certain random regular graphs. [12] does an investigation in the case of a generalized small world model. On the other hand, [26] discusses a natural generalization of the standard clustering coefficient, and analyzes in both Erdős-Rényi and small world settings.
Now, similar to the degree index, for $\alpha>0$, we define the $\alpha$-level clustering index of $\mathcal{G}$ to be
@@
Here again $\alpha$ in our case is either 1 or 2. As noted above, although such an index seems natural to define, it is first introduced in this manuscript to the best of our knowledge. Before continuing with our contributions, let us now include some discussion of this newly introduced index.
The clustering index $\mathrm{CI}_{\alpha}(\mathcal{G})$ captures the heterogeneity in local clustering across the graph, offering insights that traditional global or average measures may miss. For instance, it can reveal whether certain nodes are embedded in tightly-knit communities where others are more isolated. This distinction is crucial in real-world networks, where some nodes may form dense subgraphs, while others may lie on the periphery with sparse neighborhoods [7]. In many systems, the core nodes tend to have high clustering, whereas the peripheral nodes exhibit low clustering. A high clustering index reflects this inequality, which could remain hidden when only considering global or average clustering coefficients.
Moreover, in networks with community structure ([20]), nodes within the same community often display similar clustering levels. Hence, a low clustering index may indicate structural homogeneity, while a high value can signal the presence of multiple structurally distinct communities. As a last note, when analyzing temporal networks, the global clustering might remain stable over time, even as the roles of nodes shift. In contrast, the clustering index is sensitive to such changes in clustering variability, making it a possible tool for detecting structural transitions or evolving roles in dynamic graphs.
Moving to the contributions of this article, our primary goal is to analyze $\mathrm{DI}_{\alpha}(G)$ and $\mathrm{CI}_{\alpha}(G)$ when $G$ is an Erdős-Rényi graph. For such random graphs, the degree index case turns out to be relatively straightforward, and precise formulas for the expected degree index are obtainable. In particular, the results in Section 4 tell that when $G$ is an Erdős-Rényi graph with $n$ nodes and a fixed attachment probability $p$, we have,
$\mathbb{E}[\mathrm{DI}_{1}(\mathcal{G})]\sim\frac{2}{\sqrt{\pi}}\binom{n}{2}\sqrt{(n-2)p(1-p)}\quad\text{and}\quad\mathbb{E}[\mathrm{DI}_{2}(\mathcal{G})]=6\binom{n}{3}p(1-p).$
On the other hand, the computations for the case of the expected clustering index are more involved, and we are not able to obtain exact formulas. However, for the same random graph family, we show that the bounds
$\mathbb{E}[\mathrm{CI}_{1}(\mathcal{G})]\leq K_{1}n\quad\text{and}\quad\mathbb{E}[\mathrm{CI}_{2}(\mathcal{G})]\leq K_{2},$
hold for every $n$, for some constants $K_{1},K_{2}$ depending on $p$, but not on $n$. Our heuristic arguments yield that matching lower bounds are true for both cases, but we have not verified these rigorously yet. Moreover, we will complement our theoretical results by making observations on the two indices of interest for other random graph models via Monte Carlo simulations. The models we study include random regular graphs, the Barabási-Albert model, and the Watts-Strogatz model, and relevant simulations and discussions will be in Section 5.1.
@@
The remainder of the paper is organized as follows. The next section contains some basic observations on the clustering index, which can be easily adapted to the degree index. This section also includes a brief discussion on the comparison of the extremal cases of the degree index and the clustering index. Afterwards, in Section 3 we analyze the clustering index in Erdős-Rényi graphs, and provide upper bounds for the expected clustering index for both $\alpha=1$ and $2$. Later, in Section 4, a similar study is done for the degree index. Section 5 is then devoted to a simulation-based study for the indices under study for random regular graphs, the Barabási-Albert model, and the Watts-Strogatz model. Lastly, we conclude the paper in Section 6 with some discussions on possible future work.
@@ 2 Some basic observations on the clustering index
In this section, we make some elementary observations on the clustering index, beginning with the analysis of the extremal values. Below, whenever it is clear from the context, we assume that the node set of the underlying graph is $V=\{1,2,\ldots,n\}$ for some $n\in\mathbb{N}$. The following elementary lemma will be used to determine the extremal values of $\mathrm{CI}_{\alpha}(\mathcal{G})$.
Lemma 2.1
Let $a_{1},a_{2},\ldots,a_{n}$ be real numbers on $[0,1]$. Then for any $\alpha\geq 1$, we have,
@@
The reason for the last equality is that for the terms containing $b_{k}$, we must have $i\leq k-1$ and $j\geq k$. The number of such pairs is (k1)(n+1k)(k-1)(n+1-k). The maximum value (k1)(n+1k)(k-1)(n+1-k) can take is $\frac{n^{2}}{4}$. Therefore, $S\leq\frac{n^{2}}{4}\sum_{k=2}^{n}b_{k}=$ $\frac{n^{2}}{4}\left(a_{n}-a_{1}\right)\leq\frac{n^{2}}{4}$. $\square$
Now, the next corollary provides the extremal values of $\mathrm{CI}_{\alpha}(\mathcal{G})$. The lower bound in the corollary is clear, and the upper bound follows from Lemma 2.1.
Corollary 2.1
For any graph $\mathcal{G}=(V,E)$ and for any $\alpha\geq 1$, we have
$0\leq\mathrm{CI}_{\alpha}(\mathcal{G})\leq\frac{n^{2}}{4}.$
@@
In this example, we discuss some elementary graphs for which the bounds stated in Corollary 2.1 are attained. Let $\alpha\geq 1$ be arbitrary. The lower bound in Corollary 2.1 is attained, for example, for the following graphs:
@@
As a simple example where the clustering index is maximized, let $m\geq 2$, $n=2m$, and consider the graph that is the graph union of an $m$-null set, and $K_{m}$ (complete graph with $m$ vertices). Figure 1 exemplifies such a construction when $n=8$.
Here, the only contribution to the sum defining $\mathrm{CI}_{\alpha}$ is from the cases where a node from $K_{m}$ and another node from the $m$-null set are considered. Noting that the clustering coefficient of the node from the $m$-null set is 0 by definition, the observation that $\mathrm{CI}_{\alpha}$ takes the value $n^{2}/4$ follows immediately.
Complete Graph $K_{4}$Null Graph $N_{4}$
Figure 1: A graph with 88 vertices where $\mathrm{CI}_{\alpha}=8^{2}/4=16$.
Another related example will be in Example 2.3 where extremal situations in clustering and degree indices are compared and discussed.
Noting that the observations in Corollary 2.1 can be easily adapted to the degree index, we continue with the following result which provides a way to express the clustering index in terms of the order statistics of the corresponding local clustering coefficients.
Lemma 2.2
@@
Proof. Following the proof of Lemma 2.1 with $a_{i}=C_{(i)}$ gives
@@
Let us also examine extremal cases in random graphs. Clearly, any random tree model provides a random graph model for which the lower bound in Corollary 2.1 is achieved. The following example, which is a randomized version of the construction given in Example 2.1, provides a simple random graph model where the upper bound in Corollary 2.1 is achieved asymptotically.
@@
Begin with two empty sets of nodes $V_{1}$ and $V_{2}$ at time 0. At each time $t\in\mathbb{N}$, a newly generated node $v_{t}$ is inserted in $V_{1}$ or $V_{2}$ with probabilities $p$ and 1p1-p, respectively. If $v_{t}$ is inserted in $V_{1}$, we do not attach it to any present nodes. If it is inserted in $V_{2}$, then we attach it to all vertices present in $V_{2}$. Then, at time $n\in\mathbb{N}$, the vertices in $V_{1}$ form a null graph, and the vertices in $V_{2}$ form a complete graph.
Letting $N$ be the number of vertices in $V_{1}$ at time $n$, NN is binomially distributed with parameters $n$ and $p$, and we have,
@@
In the following last example, we provide a comparison of the clustering index and the degree index. In particular, we are interested in cases where the degree index is high and the clustering coefficient is low, and vice versa.
@@
Consider a collection of disjoint polygons, i.e. a graph where each node is contained in a unique polygon. $C(i)=1$ if $i$ is contained in a triangle, and $C(i)=0$ if $i$ is contained in a polygon other than a triangle. The degree of any node in this graph is 2. Assuming we have $k$ vertices contained in some triangle and nkn-k vertices contained in a polygon which is not a triangle, we have $\mathrm{CI}_{\alpha}=\sum\limits_{1\leq i<j\leq n}|C(i)-C(j)|^{\alpha}=nk-k^{2}$. If $n$ is even, then the maximum is attained when $k=n/2$, which yields $n^{2}/4$. Thus we have found a graph for which the degree index is zero, but the clustering index is $n^{2}/4$, the maximal possible value.
@@
Next, we also discuss an example where the clustering index is zero while the degree index $\sum\limits_{1\leq i<j\leq n}|d_{i}-d_{j}|$ is $\frac{n^{2}(n/2-3)}{4}$. (Lemma 2.1 can be used to show that the degree index can be at most $n^{3}/4$. So our example here is off by a factor of $1/2$, but still demonstrates that the degree index can be very large while the clustering index is zero.)
@@ 3 Clustering index of Erdős-Rényi graphs
In this subsection we will focus on the case where $\mathcal{G}$ is an Erdős-Rényi graph, and show that $\mathbb{E}[\mathrm{CI}_{1}(\mathcal{G})]$ and $\mathbb{E}[\mathrm{CI}_{2}(\mathcal{G})]$ are both sublinear. The result for the latter will later be improved, and in particular $\mathbb{E}[\mathrm{CI}_{2}(\mathcal{G})]$ will be shown to be bounded by a constant independent of $n$. We begin with the following proposition concerning the first two moments of the local clustering coefficient.
@@
$\mathbb{E}[C(i)]=\phi\mathbb{E}[C(i)\mid d_{i}\leq 1]+(1-\phi)\mathbb{E}[C(i)\mid d_{i}\geq 2].$ (1)
Now, by the definition of $C(i)$, $\mathbb{E}[C(i)\mid d_{i}\leq 1]=0$. We claim $\mathbb{E}[C(i)\mid d_{i}\geq 2]=p$. We have
$\mathbb{E}[C(i)\mid d_{i}\geq 2]=\mathbb{E}[\mathbb{E}[C(i)\mid d_{i},d_{i}\geq 2]\mid d_{i}\geq 2]=\mathbb{E}\left[\frac{1}{\binom{d_{i}}{2}}\mathbb{E}\left[\sum_{k,\ell\in N(i)}\mathbf{1}(k\leftrightarrow\ell)\mid d_{i},d_{i}\geq 2\right]\mid d_{i}\geq 2\right],$
where we recall that $N(i)$ is the set of neighbors of node $i$. Now given $d_{i}$ with $d_{i}\geq 2$, $\sum_{k,\ell\in N(i)}\mathbf{1}(k\leftrightarrow\ell)$ is binomially distributed with parameters $\binom{d_{i}}{2}$ and $p$, whose expectation is $\binom{d_{i}}{2}p$. Therefore, continuing the last observations, we conclude,
$\mathbb{E}[C(i)\mid d_{i}\geq 2]=\mathbb{E}\left[\frac{1}{\binom{d_{i}}{2}}\binom{d_{i}}{2}p\right]=p.$
Substituting this into (1), we obtain
$\mathbb{E}[C(i)]=p(1-(1-p)^{n-1}-(n-1)p(1-p)^{n-2})=p-p(1-p)^{n-1}-(n-1)p^{2}(1-p)^{n-2}.$
Now, for $p\in(0,1)$, the sequence $p(1-p)^{n-1}+(n-1)p^{2}(1-p)^{n-2}$ decays exponentially fast as $n\rightarrow\infty$ and thus, it is upper bounded by $D/n^{2}$ for some $D>0$. The assertion that $\mathbb{E}[C(i)]\geq p-\frac{D}{n^{2}}$ is now clear.
(ii) Note that below, $D$ denotes a constant independent of $n$ that does not necessarily have the same value in its two appearances.
Let $A_{i}$ be the event that $|d_{i}-(n-1)p|<n^{2/3}$. Then by McDiarmid (or Azuma-Hoeffding) inequality [19], $\mathbb{P}(\overline{A_{i}})\leq M_{1}e^{-M_{2}n^{1/3}}$ for some positive constants $M_{1},M_{2}$ independent of $n$. Using $A_{i}$, we now write,
@@
Now, when $A_{i}$ is true, $(n-1)p-n^{2/3}\leq d_{i}\leq(n-1)p+n^{2/3}$. Keeping this in mind, and using the last relation we obtained, we get,
@@
Proof. Let $i,j\in\mathbb{N}$. Note that $C(i)$ and $C(j)$ have the same distribution in Erdős-Rényi graphs due to the underlying symmetry. Using this along with the Cauchy-Schwarz inequality, and the previous proposition, we observe,
@@
In the next subsection, we will show that in the case of Erdős-Rényi graphs for fixed $p$, $\mathbb{E}[\mathrm{CI}_{2}(\mathcal{G})]$ is indeed bounded by a constant that does not depend on $n$. Note that our simulation results further suggest that for Erdős-Rényi graphs, the growth of $\mathrm{CI}_{1}(\mathcal{G})$ behaves linearly in the number of nodes, and $\mathrm{CI}_{2}(\mathcal{G})$ converges to a positive constant as the number of nodes increases.